OpenStudy (anonymous):

Find the equation of a line that passes through the point (4,2) that is perpendicular to the line y = 4/3 x. Show your work.

OpenStudy (anonymous):

you know the slope of \[y=\frac{4}{3}x\]?

OpenStudy (anonymous):

No, I am confused on the whole thing.

OpenStudy (anonymous):

hint it is the only number you see

OpenStudy (natsent06):

satellite after this question can u plz come look at mine

OpenStudy (anonymous):

k

OpenStudy (natsent06):

thanks

OpenStudy (anonymous):

I am confused because i need the work draw out so i can answer on paper.

OpenStudy (anonymous):

and i never worked with perpendicular piles

OpenStudy (anonymous):

forget about perpendicular for a moment step one is to find (with your eyeballs) the slope of \[y=\frac{4}{3}x\] which is pretty much the only number you are looking at

OpenStudy (anonymous):

all i was told is to flip it.

OpenStudy (anonymous):

can we take this one step at a time?

OpenStudy (anonymous):

yes but my teacher all he said was get the slope and flip it. And i cant work it out.

OpenStudy (anonymous):

forget what your teacher said it is wrong in any case we need first to find the "it" i.e. we need to find the slope of \[y=\frac{4}{3}x\] can you please write the slope?

OpenStudy (anonymous):

thats my problem, I can only do slope of two points.

OpenStudy (anonymous):

oh i see you find the slope if you are told two points by computing it but here you are not told two points, you are told the equation of the line therefore you are also told the slope it is the only number you see there

OpenStudy (anonymous):

if you have \[y=mx+b\] the slope is \(m\) and the \(y\) intercept is \(b\) you have \[y=\frac{4}{3}x\] what is \(m\) in this example?

OpenStudy (anonymous):

.... 4/3 ?

OpenStudy (anonymous):

\[\huge \text{whew!!}\]

OpenStudy (anonymous):

how would it go from there.... to make the equation.

OpenStudy (anonymous):

ok now we have step one, find the slope which really means just write it since there is nothing to find now on to step two

OpenStudy (anonymous):

for the slope of the "perpendicular line" you have to do two things take your slope (which we now know is \(\frac{4}{3}\) and a) flip it b) change the sign the slope of your perpendicular line is therefore \[-\large \frac{3}{4}\]

OpenStudy (anonymous):

so the equation would be y=-3/4x + b ?

OpenStudy (anonymous):

right that is it you need \(b\) and there are two ways to find it

OpenStudy (anonymous):

the point is \( (4,2)\) and the slope is \(-\frac{3}{4}\) use the point slope formula and write \[y-2=-\frac{3}{4}(x-4)\] and solve for \(y\) is one way

OpenStudy (anonymous):

dont you distribute the -3/4...? Cause that's where i get lost.

OpenStudy (anonymous):

yes you do

OpenStudy (anonymous):

\[y-2=-\frac{3}{4}x+3\] is the first step

OpenStudy (anonymous):

because \(-\frac{3}{4}\times -4=3\)

OpenStudy (anonymous):

last step is to add \(2\) and get \[y=-\frac{3}{4}x+5\]

OpenStudy (anonymous):

where does the minus 4 come from